# TSD-SR: One-Step Diffusion with Target Score Distillation for Real-World Image Super-Resolution

### v3

@@ Abstract
Pre-trained text-to-image diffusion models are increasingly applied to real-world image super-resolution (Real-ISR) task. Given the iterative refinement nature of diffusion models, most existing approaches are computationally expensive. While methods such as SinSR and OSEDiff have emerged to condense inference steps via distillation, their performance in image restoration or details recovery is not satisfied. To address this, we propose TSD-SR, a novel distillation framework specifically designed for real-world image super-resolution, aiming to construct an efficient and effective one-step model. We first introduce the Target Score Distillation, which leverages the priors of diffusion models and real image references to achieve more realistic image restoration. Secondly, we propose a Distribution-Aware Sampling Module to make detail-oriented gradients more readily accessible, addressing the challenge of recovering fine details. Extensive experiments demonstrate that our TSD-SR has superior restoration results (most of the metrics perform the best) and the fastest inference speed (e.g. 40 times faster than SeeSR) compared to the past Real-ISR approaches based on pre-trained diffusion priors. Our code will be released at https://github.com/Microtreei/TSD-SR.
@@ 1 Introduction
Image super-resolution (ISR) [9, 8, 23, 26] aims to transform low-quality (LQ) images, which have experienced noise or blur, into clear high-quality (HQ) images. Differing from traditional ISR [6, 67], which assumes a known degradation process, real-world image super-resolution (Real-ISR) [47, 63] is designed to enhance real-world images that have suffered from complex and unknown degradations, thereby offering greater practical utility.
Generative models, particularly Generative Adversarial Networks (GANs) [11, 33, 37] and Diffusion Models (DMs) [42, 17, 40], have shown remarkable power in handling Real-ISR task. GAN-based methods leverage adversarial training, toggling between the generator and discriminator to produce realistic images. While GANs are capable of one-step inference, they are commonly hampered by issues like mode collapse and training instability [2]. Recently, Diffusion Models (DMs) have shown impressive performance in the realm of image generation [48, 21]. Their robust priors empower them to produce more realistic images with richer details than GAN-based methods [40, 42]. Some researchers [29, 61, 54, 57] have successfully leveraged pre-trained DMs for Real-ISR task. However, due to the iterative denoising nature of diffusion models [17], the Real-ISR process is computationally expensive.
@@
To achieve an efficient and one-step network akin to GANs, several pioneering methods that condense the iterations of diffusion models through distillation [15, 58, 12, 18] have been proposed [53, 55, 49]. Among these works, OSEDiff [53] introduced the Variational Score Distillation (VSD) loss [51] to Real-ISR task, achieving state-of-the-art (SOTA) one-step performance by leveraging prior knowledge from pre-trained models. Despite these advancements, our investigation has revealed two critical limitations associated with VSD in Real-ISR applications. (1) Unreliable gradient direction. VSD relies on a Teacher Model to provide a “true gradient direction.” However, this guidance proves unreliable in scenarios where initial ISR outputs are suboptimal. (2) Insufficient detail recovery. The VSD loss exhibits notable variation across different timesteps, and the uniform sampling strategy for $t$ poses challenges in aligning the score function with detailed texture recovery requirements. These findings highlight the need for enhanced approaches to address these issues effectively.
In this paper, we propose a novel method called TSD-SR to distill a multi-step Text-to-Image (T2I) DMs [38, 40, 10] into an effective one-step diffusion model tailored for the Real-ISR task. Specifically, TSD-SR consists of two components: Target Score Distillation (TSD) and Distribution-Aware Sampling Module (DASM). TSD incorporates our new proposed Target Score Matching (TSM) loss to compensate for the limitation of VSD loss. This significant score loss leverages HQ data to provide a reliable optimization trajectory for the distillation process. It effectively reduces the visual artifacts caused by deviant predictions from the Teacher Model. DASM is crafted to bolster detail recovery by strategically sampling low-noise samples that are distribution-based during training. This approach effectively allocates more optimization to early timesteps in a single iteration, enhancing the recovery of details.
Experiments on popular benchmarks demonstrate that TSD-SR achieves superior restoration performance (most of the metrics perform the best) and high efficiency (the fastest inference speed, 40 times faster than SeeSR) compared to the state-of-the-art Real-ISR methods based on pre-trained DMs, while requiring only a single inference step.
@@
  • We introduce Target Score Distillation (TSD) to provide a reliable gradient to enhance the realism of the Real-ISR model’s outputs.
@@ 2 Related Work
One-step Diffusion-based Real-ISR. Recently, there has been a surge of interest within the academic community in one-step distillation techniques [34, 60, 41, 59, 31] for diffusion-based Real-ISR task. SinSR [49] leverages consistency preserving distillation to condense the inference steps of ResShift [62] into a single step, yet the generalization of ResShift and SinSR is constrained due to the absence of large-scale data training. AddSR [55] introduces the adversarial diffusion distillation (ADD) [41] to Real-ISR tasks, resulting in a comparatively effective four-step model. However, this method has a propensity to produce excessive and unnatural image details. OSEDiff [53] directly uses LQ images as the beginning of the diffusion process, and employs VSD loss [51] as a regularization technique to condense a multi-step pre-trained T2I model into a one-step Real-ISR model. However, due to the incorporation of alternating training strategies, OSEDiff may initially tend towards unreliable optimization directions, which may lead to the visual artifacts.
@@ 3 Methodology
Figure 2: Pipeline overview. We train one-step Student Model $G_{\theta}$ to map the low-quality image $x_{L}$ into a more realistic one. The noisy latent $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ sampled by DASM (Details can be found in Fig. 6.) will be fed into both the pre-trained Teacher and LoRA Model to produce the Variational Score loss. Subsequently, two Models’ predictions on $\boldsymbol{z_{t}}$ yield Target Score loss. Their weighted forms, namely TSD ( red flow), in conjunction with pixel-space reconstruction loss ( green flow), are leveraged to update the Student Model. After updating the Student Model, we employ the diffusion loss ( blue flow) to alternately update the LoRA Model.
@@ 3.1 Preliminaries
Problem Formulation. ISR problem aims to reconstruct a HQ image ${x}_{H}$ from a LQ input $x_{L}$ by training a parameterized ISR model $G_{\theta}$ on a dataset $\mathcal{D}=\{(x_{L},x_{H})_{i=1}^{N}\}$, where $N$ represents the number of image pairs. It is mathematically formulated as minimizing the following objective:
@@
Here, $\mathcal{L}_{Rec}$ denotes the reconstruction loss, commonly measured using distance metrics like $L_{2}$ or $LPIPS$ [66]. The regularization term $\mathcal{L}_{Reg}$ enhances the realism and generalization of the ISR model’s outputs. This objective can be understood as aligning the ISR output $\hat{x}_{H}$’s distribution, $q_{\theta}(\hat{x}_{H})$, with the true distribution $p(x_{H})$ by minimizing the KL-divergence [25]:
@@
While several studies [47, 63, 46] have employed adversarial loss to optimize this objective, they often encounter issues like mode collapse and training instability. Recent work [53] achieved the state-of-the-art results using Variational Score Distillation (VSD) as the regularization loss to minimizing this objective, which inspires our research.
Variational Score Distillation. Variational Score Distillation (VSD) [51] was initially introduced for text-to-3D generation by distilling a pre-trained text-to-image diffusion model to optimize a single 3D representation [36].
@@ 3.2 Overview of TSD-SR
As depicted in Fig. 2, our goal is to distill a given pre-trained T2I DMs into a fast one-step Student Model $G_{\theta}$ with Teacher Model $\boldsymbol{\epsilon}_{\psi}$ and trainable LoRA Model $\boldsymbol{\epsilon}_{\phi}$. We denote the latent outputs of the distilled model as $\boldsymbol{\hat{z}_{0}}$, and HQ latent representations as $\boldsymbol{z_{0}}$. Both $\boldsymbol{\hat{z}_{0}}$ and $\boldsymbol{z_{0}}$ will pass through our Distribution-Aware Sampling Module (DASM) to obtain distribution-based samples $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ (Sec. 3.4). We train $G_{\theta}$ by minimizing the two losses: a reconstruction loss in pixel space to compare the model outputs against the ground truth, and a regularization loss (from Target Score Distillation) to enhances the realism (Sec. 3.3). After updating the Student Model, we update the LoRA Model with diffusion loss. Finally, in Sec. 3.5, we present an overview of all the losses encountered during the training phase.
@@ 3.3 Target Score Distillation
Similar to [53], we introduce VSD loss into our work as a regularization loss to enhance the realism and generalization of the $G_{\theta}$ outputs. Upon reviewing VSD Eq. 3, $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ represents the gradient direction of the optimization on $G_{\theta}$ noisy outputs $\boldsymbol{\hat{z}_{t}}$, whereas $\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ corresponds to the gradient direction on noisy HQ latent representations $\boldsymbol{z_{t}}$. The overarching goal of model optimization is to harmonize the predictions from $\boldsymbol{\hat{z}_{t}}$ with those from $\boldsymbol{z_{t}}$. Nonetheless, this strategy encounters hurdles, especially in the early training phase: the quality of synthetic latents $\boldsymbol{\hat{z}_{t}}$ is not high enough for Teacher Model to give a precise prediction. As illustrated in Fig. 3, the Teacher Model finds it challenging to accurately predict the optimization direction for low-quality synthetic latents, with the cosine similarity to the ideal optimization direction only $0.2$ ($0.88$ on HQ latents). This problem can lead to severe visual artifacts, as evident in Fig. 4(a).
Figure 3: A visual comparison of the gradient direction. We set the timestep $t$ to 100 and calculated the cosine similarity between the predicted directions from the Teacher Model and the true direction (towards the HQ data). The prediction direction for $\boldsymbol{z_{t}}$ closely matches the true direction, but not for $\boldsymbol{\hat{z}_{t}}$, suggesting that suboptimal samples may lead to directional deviations.
@@
Figure 4: The visualization of different strategies. (a) The naive method introduces fake textures and fails to recover fine details. (b) MSE leads to over-smoothed generation result, lacking high-frequency information. (c) Our method offers the superior visual effects and fine textures.
A straightforward remedial measure is to employ a mean squared error (MSE) loss to align the synthetic latents with the ideal inputs of the Teacher Model, which are derived from HQ latents. However, as shown in Fig. 4(b), this approach has been observed to lead to over-smoothed results [13]. Our strategy, instead, is to align the predictions made by the Teacher Model on both synthetic and HQ latents, thereby encouraging greater consistency between them. The core idea is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. We refer to this approach as Target Score Matching (TSM):
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\right]\end{split}$ (4)
where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$. Equation 4 encapsulates the optimization loss for our Target Score Matching. Upon examining it in conjunction with the Eq. 3, we notice that VSD utilizes the prediction residual between a Teacher and a LoRA Model to drive gradient backpropagation. Conversely, our TSM employs the synthetic and the HQ data to produce the gradients. By blending these two strategies with hyperparameter weighting $\lambda$ and $1-\lambda$, we can construct our ultimate optimization loss, effectively unifying the strengths of both approaches to optimize the training process Eq. 5.
$\begin{split}\nabla_{\theta}\mathcal{L}_{\mathrm{TSD}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)[{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\\ {\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})+\lambda({\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})-{\epsilon}_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y}))]\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\Big{]}\end{split}$ (5)
where $w(t)$ is time-aware weighting function tailored for Real-ISR. Other symbols are in accordance with those previously mentioned. By introducing the prediction of pre-trained diffusion model on HQ latents, we have circumvented the issue of the model falling into the visual artifacts or over-smoothed problem, as illustrated in Fig. 4(c).

### v4

@@ Abstract
Pre-trained text-to-image diffusion models are increasingly applied to real-world image super-resolution (Real-ISR) tasks. Given the iterative refinement nature of diffusion models, most existing approaches are computationally expensive. While methods such as SinSR and OSEDiff have emerged to condense inference steps via distillation, their performance in image restoration or details recovery is not satisfactory. To address this, we propose TSD-SR, a novel distillation framework specifically designed for real-world image super-resolution, aiming to construct an efficient and effective one-step model. We first introduce the Target Score Distillation, which leverages the priors of diffusion models and real image references to achieve more realistic image restoration. Secondly, we propose a Distribution-Aware Sampling Module to make detail-oriented gradients more readily accessible, addressing the challenge of recovering fine details. Extensive experiments demonstrate that our TSD-SR has superior restoration results (most of the metrics perform the best) and the fastest inference speed (e.g. 40 times faster than SeeSR) compared to the past Real-ISR approaches based on pre-trained diffusion priors. Our code is released at https://github.com/Microtreei/TSD-SR.
@@ 1 Introduction
Image super-resolution (ISR) [9, 8, 23, 26] aims to transform low-quality (LQ) images, which have been degraded by noise or blur, into clear high-quality (HQ) images. Unlike traditional ISR methods [6, 67], which assume a known degradation process, real-world image super-resolution (Real-ISR) [47, 63] focuses on enhancing images affected by complex and unknown degradations, thereby offering greater practical utility.
Generative models, particularly Generative Adversarial Networks (GANs) [11, 33, 37] and Diffusion Models (DMs) [42, 17, 40], have demonstrated remarkable capabilities in tackling Real-ISR tasks. GAN-based methods utilize adversarial training by alternately optimizing a generator and a discriminator to produce realistic images. While GANs support one-step inference, they are often hindered by challenges such as mode collapse and training instability [2]. Recently, Diffusion Models (DMs) have demonstrated impressive performance in image generation [48, 21]. Their strong priors enable them to produce more realistic images with richer details compared to GAN-based methods [40, 42]. Some researchers [29, 61, 54, 57] have successfully leveraged pre-trained DMs for Real-ISR tasks. However, due to the iterative denoising nature of diffusion models [17], the Real-ISR process is computationally expensive.
@@
To achieve an efficient and one-step network akin to GANs, several pioneering methods that condense the iterations of diffusion models through distillation [15, 58, 12, 18] have been proposed [53, 55, 49]. Among these works, OSEDiff [53] introduced the Variational Score Distillation (VSD) loss [51] to Real-ISR tasks, achieving state-of-the-art (SOTA) one-step performance by leveraging prior knowledge from pre-trained models. Despite these advancements, our investigation has revealed two critical limitations associated with VSD in Real-ISR applications. (1) Unreliable gradient direction. VSD relies on a Teacher Model to provide a “true gradient direction.” However, this guidance is proven unreliable in scenarios where initial ISR outputs are suboptimal. (2) Insufficient detail recovery. The VSD loss exhibits notable variation across different timesteps, and the uniform sampling strategy for $t$ poses challenges in aligning the score function with detailed texture recovery requirements. These findings underscore the need for more effective approaches to address these issues.
In this paper, we propose a novel method called TSD-SR to distill a multi-step Text-to-Image (T2I) DMs [38, 40, 10] into an effective one-step diffusion model tailored for the Real-ISR task. Specifically, TSD-SR consists of two components: Target Score Distillation (TSD) and Distribution-Aware Sampling Module (DASM). TSD incorporates our newly proposed Target Score Matching (TSM) loss to compensate for the limitations of the VSD loss. This significant score loss leverages HQ data to provide a reliable optimization trajectory during distillation, effectively reducing visual artifacts caused by deviant predictions from the Teacher Model. DASM is designed to enhance detail recovery by strategically sampling low-noise samples that are distribution-based during training. This approach effectively allocates more optimization to early timesteps within a single iteration, thereby improving the recovery of fine details.
Experiments on popular benchmarks demonstrate that TSD-SR achieves superior restoration performance (most of the metrics perform the best) and high efficiency (the fastest inference speed, 40 times faster than SeeSR) compared to the state-of-the-art Real-ISR methods based on pre-trained DMs, while requiring only a single inference step, as shown in (Fig. 1).
@@
  • We introduce Target Score Distillation (TSD) to provide reliable gradients that enhance the realism of outputs from Real-ISR methods.
@@ 2 Related Work
One-step Diffusion-based Real-ISR. Recently, there has been a surge of interest within the academic community in one-step distillation techniques [34, 60, 41, 59, 31] for diffusion-based Real-ISR task. SinSR [49] leverages consistency preserving distillation to condense the inference steps of ResShift [62] into a single step, yet the generalization of ResShift and SinSR is constrained due to the absence of large-scale data training. AddSR [55] introduces the adversarial diffusion distillation (ADD) [41] to Real-ISR tasks, resulting in a comparatively effective four-step model. However, this method has a propensity to produce excessive and unnatural image details. OSEDiff [53] directly uses LQ images as the beginning of the diffusion process, and employs VSD loss [51] as a regularization technique to condense a multi-step pre-trained T2I model into a one-step Real-ISR model. However, due to the incorporation of alternating training strategies, OSEDiff may initially tend towards unreliable optimization directions, which may lead to visual artifacts.
@@ 3 Methodology
Figure 2: Pipeline overview. We train a one-step Student Model $G_{\theta}$ to transform the low-quality image $x_{L}$ into a more realistic one. The noisy latent $\boldsymbol{\hat{z}_{t}}$ sampled by DASM (Details can be found in Fig. 6.) will be fed into both the pre-trained Teacher and the LoRA Model to produce the Variational Score Loss. Subsequently, the Teacher’s predictions on $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ yield the Target Score Loss. Their weighted forms, namely TSD ( red flow), along with the pixel-space reconstruction loss ( green flow), are leveraged to update the Student Model $G_{\theta}$ After updating the Student Model, we employ the diffusion loss ( blue flow) to update the LoRA Model.
@@ 3.1 Preliminaries
Problem Formulation. The ISR problem aims to reconstruct a HQ image ${x}_{H}$ from an LQ input $x_{L}$ by training a parameterized ISR model $G_{\theta}$ on a dataset $\mathcal{D}=\{(x_{L},x_{H})_{i=1}^{N}\}$, where $N$ represents the number of image pairs. Formally, this problem can be formulated as minimizing the following objective:
@@
Here, $\mathcal{L}_{Rec}$ denotes the reconstruction loss, commonly measured by distance metrics such as $L_{2}$ or $LPIPS$ [66]. The regularization term $\mathcal{L}_{Reg}$ improves the realism and generalization of the output of the ISR model. This objective can be understood as aligning the ISR output $\hat{x}_{H}$’s distribution, $q_{\theta}(\hat{x}_{H})$, with the high-quality data $x_{H}$’s distribution $p(x_{H})$ by minimizing the KL-divergence [25]:
@@
While several studies [47, 63, 46] have employed adversarial loss to optimize this objective, they often encounter issues like mode collapse and training instability. Recent work [53] achieved state-of-the-art results using Variational Score Distillation (VSD) as the regularization loss to minimize this objective, which inspires our research.
Variational Score Distillation. Variational Score Distillation (VSD) [51] was initially introduced for text-to-3D generation, by distilling a pre-trained text-to-image diffusion model to optimize a single 3D representation [36].
@@ 3.2 Overview of TSD-SR
As depicted in Fig. 2, our goal is to distill a given pre-trained T2I DM into a fast one-step Student Model $G_{\theta}$, using the Teacher Model $\boldsymbol{\epsilon}_{\psi}$ and the trainable LoRA Model $\boldsymbol{\epsilon}_{\phi}$. We denote the latent output of the distilled model as $\boldsymbol{\hat{z}_{0}}$, and the HQ latent representation as $\boldsymbol{z_{0}}$. Both $\boldsymbol{\hat{z}_{0}}$ and $\boldsymbol{z_{0}}$ are passed through our Distribution-Aware Sampling Module (DASM) to obtain distribution-based samples $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ (Sec. 3.4). We train $G_{\theta}$ by minimizing the two losses: a reconstruction loss in pixel space to compare the model outputs against the ground truth, and a regularization loss (from Target Score Distillation) to enhance the realism (Sec. 3.3). After updating the Student Model, we update the LoRA Model with the diffusion loss. Finally, in Sec. 3.5, we present an overview of all the losses encountered during the training phase.
@@ 3.3 Target Score Distillation
Similar to [53], we introduce VSD loss into our work as a regularization term to enhance the realism and generalization of the $G_{\theta}$’s outputs. Upon reviewing VSD Eq. 3, $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ represents the current estimated gradient direction for $G_{\theta}$’s noisy outputs $\boldsymbol{\hat{z}_{t}}$, whereas $\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ corresponds to the ideal gradient direction guiding towards more realistic outputs. The overarching goal of model optimization is to align the suboptimal gradient direction with the superior direction based on pre-trained priors, thus facilitating the optimization of the Student distribution toward that of the Teacher. However, this strategy encounters hurdles, especially in the early training phase: the quality of synthetic latent $\boldsymbol{\hat{z}_{t}}$ is not high enough for the Teacher Model to provide a precise prediction. As illustrated in Fig. 3, the Teacher Model struggles to accurately predict the optimization direction for low-quality synthetic latent $\boldsymbol{\hat{z}_{t}}$ in the early stage, as indicated by a cosine similarity of only $0.2$ to the ideal direction, compared to $0.88$ for high-quality latent $\boldsymbol{z_{t}}$. This problem can lead to severe visual artifacts, as is evident in Fig. 4(a).
Figure 3: A visual comparison of the gradient direction. We set the timestep $t$ to 100 and calculated the cosine similarity between the prediction directions from the Teacher Model and the true direction (towards the HQ data). The prediction direction for $\boldsymbol{z_{t}}$ closely matches the true direction, but not for $\boldsymbol{\hat{z}_{t}}$, suggesting that suboptimal samples may lead to directional deviations.
@@
Figure 4: The visualization of different strategies. (a) The naive method introduces fake textures and fails to recover fine details. (b) MSE leads to over-smoothed generation results, lacking high-frequency information. (c) Our method offers the superior visual effects and fine textures.
A straightforward remedial measure is to employ a mean squared error (MSE) loss to align the synthetic latent with the ideal inputs of the Teacher Model, which are derived from the HQ latent. However, as shown in Fig. 4(b), this approach has been observed to lead to over-smoothed results [13]. Our strategy, instead, is to align the predictions made by the Teacher Model on both synthetic and HQ latent, thereby encouraging greater consistency between them. The core idea is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. We refer to this approach as Target Score Matching (TSM):
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y}))\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}\right]\end{split}$ (4)
where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$. Equation 4 encapsulates the optimization loss for our Target Score Matching. Upon examining it in conjunction with Eq. 3, we notice that VSD utilizes the prediction residual between the Teacher and the LoRA Model to drive gradient backpropagation. Similarly, our TSM employs the synthetic and the HQ data to produce the gradients. By blending these two strategies with hyperparameter weights $\lambda$ and $1-\lambda$, we construct a combined optimization loss that effectively unifies the strengths of both approaches, as formulated in Eq. 5, to guide the training process.
$\begin{split}\nabla_{\theta}\mathcal{L}_{\mathrm{TSD}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)[{\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-\\ {\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y})+\lambda({\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y})-{\epsilon}_{\phi}(\boldsymbol{\hat{z}}_{t};t,c_{y}))]\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}\Big{]}\end{split}$ (5)
where $w(t)$ is a time-aware weighting function tailored for Real-ISR. Other symbols are consistent with those previously defined. By introducing the prediction of the pre-trained diffusion model on HQ latent, we have circumvented the issue of the model falling into visual artifacts or producing over-smoothed results, as illustrated in Fig. 4(c).